OpenStudy (anonymous):

solve -7(11-4w)-15w=-27-(50-13w)

OpenStudy (anonymous):

-77+28w-15w=-27-50+13w 7w-77=13w-77 6w=0 w=0

OpenStudy (anonymous):

does that match with your answer